username193:

Part A: A mechanic buys 2 barrels of oil. If there are 31.5 gallons in every barrel and approximately 3.785 liters for every gallon, approximately how many liters of oil did the mechanic buy? Part B: The mechanic plans to pour the oil into 2-liter containers. How many containers will the mechanic need to hold all of the oil? Explain your answer.

username193:

could someone help me?

Matchatea:

A: 238.455 L 31.5 × 2 = 63 gallons 63 gallons = 3.785 × 63 = 238.455 L B: 120 containers 238.455 ÷ 2 = 119.0.455 119 + 1 = 120 containers.

jhonyy9:

@Matchatea nice job - congrats ! but please try explain to asker what you ve used there step by step so to make understanding all your work easy